OpenStudy (anonymous):

x+y=8 2x+7y=31 solve the system algebraically. what is the solution?

OpenStudy (anonymous):

x = 8 - y (equation 1) sub equation 1 into 2x + 7y = 31 2(8-y) + 7y = 31 16 - 2y + 7y = 31 16 + 5y = 31 5y = 31- 16 y = 15 / 5 y = 3 (equation 2) sub equation 2 into equation 1 x = 8 - y = 8 - 3 = 5